there are four more quarters than dimes there total is $2.05 how
Answer by stanbon(75887)

there are four more quarters than dimes there total is $2.05 how
-----
Equations:
q = d+4
10d + 25q = 205 cents
-----
Substitute for "q" and solve for "d":
10d + 25(d+4) = 205
35d + 100 = 205
35d = 105
d = 3 (# of dimes)
-----
Since q = d+4, q = 7 (# of quarters)
======================================
Cheers,
Stan H.
==========